OpenStudy (anonymous):

Cos^-1(-1/2)+Cot^-1(sqrt3/3)??? help pleasee

OpenStudy (anonymous):

i hate it when chrome bombs!!!

OpenStudy (anonymous):

Do you know what \(\large cos^{-1}(\frac{-1}{2}) \) means?

OpenStudy (anonymous):

isnt the same as 2pi/3?

OpenStudy (anonymous):

yes....

OpenStudy (anonymous):

how about the second term???

OpenStudy (anonymous):

no i do not

OpenStudy (anonymous):

\(\large y=cot^{-1}(\frac{\sqrt3}{3}) \rightarrow coty=\frac{\sqrt3}{3} \)... Well... let's simplify that....

OpenStudy (anonymous):

pi/4?

OpenStudy (anonymous):

\(\large coty=\frac{\sqrt3}{3} \rightarrow tany=\frac{3}{\sqrt3}=\sqrt3 \)

OpenStudy (anonymous):

do you use the unit circle?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok... tangent of what angle is square root of 3 ???

OpenStudy (anonymous):

-60

OpenStudy (anonymous):

noooo 120

OpenStudy (anonymous):

i got 60 degrees... (or pi/3)...

OpenStudy (anonymous):

nvm pi/3 is what my partner got

OpenStudy (anonymous):

ok.. so add them up and you'll get your answer.... :)

OpenStudy (anonymous):

thank you :)

OpenStudy (anonymous):

yw...:)